1. 4.990 grams of C2H4 is reacted with 31.46 gram F2 according to the following equation....

1.

4.990 grams of C2H4 is reacted with 31.46 gram F2 according to the following equation. How many gram of HF can be formed, assuming 100% yield?
C2H4 +   6F2   -->   2CF4 + 4 HF
Hint: Find limiting reagent.

6.281 g                                

8.830 g                                

4.297 g

17.21 g

20.55 g

18.34 g

11.04 g

25.91 g

2.

49.7 mL of a solution of NaOH is needed to neutralize a solution that contains 1.86 g of H2SO4  in water.    What is the molarity of the NaOH solution?

Hint--need to change g sulfuric acid to moles
  
H2SO4 + 2 NaOH   --> 2H2O + Na2SO4

0.763 M

0.0404 M   

1.13 M

0.481 M

0.602  M

2.23 M

1.81  M

2.94  M

3.

We find that 18.90 mL of a 2.50M KOH solution are required to titrate 35.0 mL of a H2SO4 solution. What is the molarity of the H2SO4 solution?
2KOH + H2SO4   --> K2SO4   + 2 H2O

0.675 M

1.57 M

0.0711  M

0.128 M

4.04 M

0.903 M

1.18 M

0.0173 M

4.

How many mL of a 1.20 M  H3PO4 solution are required to titrate 85.5 mL of a 0.465 M LiOH   solution?
H3PO4(aq) + 3LiOH (aq) --> Li3PO4(aq) +   3H2O(l).

11.0 mL

37.0 mL

6.85 mL

15.3 mL

25.1 mL

8.23 mL

5.21 mL                             

17.9 mL

5. In the following, what was reduced?  
2Na(s) + 2H2O(l)   -->   2NaOH(aq) +   H2(g)

H2

O in NaOH

H in H2O

Na

Na in NaOH
Answer #1

Question 1

Balanced equation
C2H4 + 6 F2 ====> 2 CF4 + 4 HF

Mass of C2H4 = 4.99 gm

Moles of C2H4 = 4.99 g / 28.05 g/mol = 0.1778 Moles

Mass of F2 =  31.46 gm

Moles of F2 = 31.46 g / 37.99 g/mol = 0.8279 Moles

Limiting reagent is F2  

Moles of HF formed =  0.8279 Moles x 4/6 =  0.5519 Moles

Mass of HF produced =  0.5519 mol x 20.00 g/mol =  11.04 g

Option G is correct answer

Question 2  

H2SO4 + 2 NaOH ===> 2 H2O + Na2SO4  

Mass of H2SO4 = 1.86 gm

Moles of H2SO4 = 1.86 g / 98.047 g/mol = 0.01896 Moles

Moles of NaOH required =  0.03792 Moles

Volume of NaOH = 49.7 ml

Concentration of NaOH =  0.03792 x 1000 /49.7 = 0.7629 M

Option A is correct answer

Question 3

2 KOH + H2SO4 ===>  K2SO4 + 2 H2O

Volume of KOH = 18.9 ml

Concentration of KOH = 2.5 M

Moles of KOH = 18.9 x 2.5 / 1000  = 0.04725 Moles

Moles of H2SO4 required =  0.02362 Moles

Volume of H2SO4 = 35 ml

Concentration of H2SO4 =  0.02362 x 1000 /35 = 0.675 M

Option A is correct answer

Question 4

H3PO4(aq) + 3 LiOH(aq) ====> Li3PO4(aq) + 3 H2O(l)

Volume of LiOH = 85.5 ml

Concentration of LiOH = 0.465 M

Moles of LiOH = 85.5 x 0.465 / 1000  = 0.03975 Moles

Moles of H3PO4 required =  0.01325 Moles

Concentration of H3PO4 = 1.2  M

Volume of H3PO4 = 0.01325 x 1000 / 1.2 = 11.04 ml

Option A is correct answer

Question 5

H in H2O is reduced from +1 oxidation state to 0 oxidation state. Hence option C is correct answer